# and quantum information science

### v2

@@ 7 Quantum complexity and space-time: a more concerted approach
Connections between geometry and entanglement were first proposed by Ryu and Takayangi [5] and were made covariant by Hubeny, Rangamani and Takayanagi [237]. These authors showed that the entanglement entropies of subregions of the theory dual to an AdS spacetime are geometrized by the areas of certain extremal surfaces of the dual geometry. These developments led to major insights into the holographic duality, for example elucidating constraints on the entanglement patterns in quantum states that admit gravitational dual descriptions [238]. An analogy between the geometry of AdS spacetimes dual to conformal field theories and the graphical structure of tensor networks for constructing quantum states also provided an impetus, offering a schematic understanding of the hierarchical pattern of short and long range entanglement in holographic states [239]. Such developments motivated an important proposal by van Raamsdonk that spacetime is actually knitted together by the underlying patterns of quantum entanglement [240]. Subsequent developments have illustrated many constraints on quantum states with holographic duals including new inequalities [238] and properties such a quantum error correction [241] (see [242] for a review) that have a bearing on the complexity of constructing such states from a simple reference.
@@
Others proceeded to study the dynamics of thermalization via the AdS/CFT correspondence, especially considering the evolution of entanglement and the corresponding dual geometric quantities after quantum quenches [147, 148, 149, 243, 150, 244]. In all these cases, an initial rise of entanglement entropy ends in saturation after a thermalization time. These studies also led to an examination of the time evolution of entanglement entropy in thermo-field double (TFD) states dual to eternal black holes, i.e., to ERBs between two asymptotically AdS universes [245, 246]. The result, computed in terms of certain extremal surfaces passing through the ERB, shows that entanglement entropy grows with time in a TFD state until it saturates at equilibrium [245]. However, the black hole interior keeps growing long after the entanglement saturates. Therefore, Susskind suggested that the growth of the black hole’s interior should be encoded in the AdS/CFT correspondence by a dual quantum mechanical quantity that increases for a time exponential in the system’s size. What quantity could conceivably keep increasing after a system reaches thermal equilibrium? It is known in computer science that some computations carried out by the action of local operators can require exponentially long time or exponentially many actions of the operators, quantified as the “complexity” of the resulting state. If we think of time evolution as a computation performed by the action of a local Hamiltonian, one might guess that the complexity of the evolving quantum state, quantified in some way, may be related to the size of the black hole’s interior [247].
@@
As we discussed in Secs. 4, 5, and 6, circuit complexity, Nielsen’s complexity, and Krylov/spread complexity all show a long period of linear growth for chaotic systems and also display the switchback effect (see discussion in Sec. 6.3.2). Therefore, any dual geometric object must also show these dynamics. Consistently with this, the initial conjecture was that the volume of the ERB wormhole is dual to some notion of time-evolving complexity of the dual Thermo-field Double state [247]. However, as we discuss in Sections 7.1 and 7.2, a number of different geometric quantities all reproduce linear growth and the switchback effect [25, 26, 27, 28, 29, 30]. We also discuss a proposal in the case of two-dimensional gravity for a non-perturbative definition of the size of the ERB wormhole [31] that saturates at times exponential in the black hole entropy, reproducing the expected late time saturation of complexity (Secs. 4,5 and 6). Strictly speaking, the various candidate holographic duals of complexity described in Sec. 7.1 all have divergences that require regulation. We understand these divergences as arising in the dual CFT from the continuum nature of the theory, just like the divergences of entanglement entropy. We explain these considerations in Sec. 7.3. Proposed extensions of the ideas to subregions, geometries with defects, and black holes in de Sitter space are summarized Sec. 7.4.
@@
All the geometric quantities proposed as duals to some notion of complexity in Sec. 7.1 are diffeomorphism-invariant, and hence we expect them to map onto precise gauge-invariant quantities in the dual theory. Is there a precise map to some particular notion of complexity? In other words, can we go beyond a qualitative agreement with the linear growth and switchback effect seen in any reasonable definition of complexity, to write an equation between gravitational and CFT quantities? Very recently, precise relations have been found between the size of an Einstein-Rosen Bridge in two-dimensional Jackiw-Teitelboim gravity and the spread complexity of a dual double-scaled SYK model, matching both the linear early time classical growth and the expected late time quantum saturation [32, 248, 34, 249, 35]. There are also equalities relating the rate of change of Krylov/spread complexity and the momentum of particles in AdS spacetime. These results are reviewed in Sec. 7.5. Sec. 7.6 discusses some ideas for relating Nielsen’s complexity to dual geometries. Despite these findings, it is difficult to match geometrical observables in the gravitational setting with quantum complexity in field theory. This difficulty led to questions about the foundations of quantum computation in the context of the AdS/CFT correspondence. The existence of pseudorandom states, discussed in Sec. 7.7, can be used to argue that some entries in the holographic dictionary will be hard to compute, or alternatively, that quantum gravity efficiently solves problems that will require exponentially long times for a quantum computer. This unexpected line of research was revealed through the study of quantum complexity.
@@ 7.1 Holographic complexity proposals
Figure 10: Geometric objects defining holographic complexity proposals drawn inside the Penrose diagrams of the Schwarzschild-AdS black hole. Pictures adapted from [27, 30]. (a) Maximal-volume codimension-one surface used in the CV proposal [Eq. (112)]. (b) WDW patch used to compute CA and CV2.0 in Eqs. (113) and (114). (c) Bulk region $\mathcal{M}$ with codimension-one boundaries $\Sigma_{\pm}$ defined by extremizing Eq. (115) in the CAny proposals.
The complexity equals volume (CV) proposal [25, 6] declares that holographic complexity is proportional to the maximal volume $\mathcal{V}$ of a codimension-one spacelike surface $\mathcal{B}$ anchored at the $(d-1)$-dimensional boundary slice $\Sigma_{\rm CFT}$ where a CFT state is defined (see fig. 10(a)), i.e.,  [^27] In the context of CV conjecture, Ref. [250] proposed to measure this geometric observable in terms of the maximal time from the horizon to the final slice (approached in the late stages of evolution). This notion also shares many similarities with the standard CV proposal, and at the same time removes certain ambiguities between small and large black holes in AdS spacetime.
@@
Other holographic proposals involve gravitational observables evaluated on the Wheeler-De Witt (WDW) patch, i.e., the bulk domain of dependence of a spacelike surface anchored at the boundary slice $\Sigma_{\rm CFT}$ (see fig. 10(b)). The first conjecture, complexity equals action (CA), associates complexity with the on-shell gravitational action $I_{\rm WDW}$ [26, 27]
@@
The gravitational action receives contributions from the null boundaries of the WDW patch, as required to make the variational principle in general relativity well-defined (see [251] or appendix A of [252]). The second proposal, complexity=volume 2.0 (CV2.0), identifies complexity as the spacetime volume of the WDW patch [28]
@@
Each of these proposals presents ambiguities. CV and CV2.0 require a length scale $\ell_{\rm bulk}$ in the holographic dictionary. The CA proposal contains contributions from the null boundaries of the WDW patch, whose normal vectors depend on an arbitrary normalization constant. The last dependence can be removed by including certain counterterms in the gravitational action, but those in turn come equipped with their own (arbitrary) length scale, which we denote $L_{\rm ct}$ [251]. When studying black holes, the CA and CV2.0 proposals probe regions near the black hole singularity. A delicate cancellation between a diverging metric factor and the vanishing volume of the sphere in the angular directions near the singularity leads to a result which does not diverge. Nevertheless, the validity of the formula in the semiclassical approximation is uncertain since the region near the singularity is not well understood in this limit.
The above ambiguities were taken as features of the holographic complexity proposal, rather than bugs, since computational complexity is also subject to several ambiguities: for instance, the choice of allowed gates in a circuit, the tolerance to reach a target state, the cost of gates etc., see Sec. 4. This perspective was taken a step further with a recent proposal that went under the name complexity equals anything (CAny), i.e., complexity may be dual to any one out of an infinite class of geometric observables. CAny requires a codimension-zero bulk region $\mathcal{M}$ with future (past) boundary $\Sigma_{+}(\Sigma_{-})$ anchored at the CFT slice $\partial\Sigma_{\pm}=\Sigma_{\rm CFT}$. One defines the following functional (see fig. 10(c)) [29, 30, 47]
@@
where the shape of the boundaries $\Sigma_{\pm}$ is varied. When more than one solution exists, the one with maximal $W_{G_{2},F_{2,\pm}}$ is chosen. Once the unique codimension-zero region (¯,Σ¯±)¯subscript¯$(\bar{\mathcal{M}},\bar{\Sigma}_{\pm})$ solving Eq. (116) is identified, holographic complexity is computed by the observable
@@ 7.2 Time-dependent properties of the holographic proposals
[…]
@@
Since $v$ is cyclic for $\mathcal{L}_{\varepsilon}$ in Eq. (120), [^28] This means that $\mathcal{L}_{\varepsilon}$ depends on $\dot{v}\coloneqq dv/d\sigma$, but not on $v$. the conjugate momenta $P_{v}^{\varepsilon}=\partial\mathcal{L}_{\varepsilon}/\partial\dot{v}$ are conserved along the profiles of $\Sigma_{\varepsilon}$. The late time regime $t\rightarrow\infty$ is achieved when each of the codimension-one surfaces $\Sigma_{\pm}$ hug a final slice at constant radial coordinate, such that the momenta approach constant values, up to exponentially-suppressed terms. This leads to the linear growth [29, 30]
@@
Let us consider the switchback effect, see e.g., [25, 253, 54]. We modify the thermofield-double state as follows [254, 255]
@@
where $W_{L}$ are low-energy perturbations acting on the left boundary. These perturbations carry energies of the order of the black hole’s temperature (thermal-scale perturbations), which is assumed to be much smaller than the black hole’s mass. Let us assume that $|t_{k+1}-t_{k}|>t_{*}$, where $t_{*}$ is the scrambling time, and the label $k=0,1,\dots,n+1$ in the inequalities includes the boundary times via the definitions $t_{0}=-t_{R}$ and $t_{n+1}=t_{L}$. We also assume that, in the sequence of times $t_{0},\ldots t_{n+1}$ there are $n_{\rm sb}$ switchbacks (also called time folds): in other words, the arguments of the absolute values $|t_{k+1}-t_{k}|$ change sign $n_{\rm sb}\leq n$ times. To keep the bulk solutions as simple as possible, we assume that the perturbations are approximately spherically symmetric. The bulk geometry dual to the perturbed state (122) can be modeled by a black hole with $n\geq 1$ shock waves produced by null matter. In this setting, any CAny observable in the late time limit reads [30]
@@
The contribution $-2n_{\rm sb}t_{*}$ reflects the spreading of the perturbations into the system during a time scale measured by the scrambling time $t_{*}=\frac{1}{2\pi T}\log\left(M/E\right)\approx\frac{1}{2\pi T}\log S$, where $E\ll M$ is the energy of the shock, and $S$ the black hole’s entropy. This characterizes the switchback effect for the CAny observables. At late boundary times $(t_{L},t_{R})$, holographic complexity will grow linearly with the sum of the boundary times $t_{L}+t_{R}$, but with a delay induced by the scrambling time of the perturbation. This phenomenon was already observed in earlier studies of the CV, CV2.0 and CA proposals [25, 27, 28]. This is not a surprise. As we said, those proposals can be obtained from CAny by appropriate limiting procedures.
@@
Let us discuss other refined properties of the holographic conjectures, associated with early and intermediate times compared to the scrambling time and the inverse temperature scale. It was originally speculated that the black hole mass $M$ should provide an upper bound on the complexity rate $d\mathcal{C}/dt\leq 2M/\pi$, based on Lloyd’s bound on the computational speed [256]. This statement was supported by the the full time evolution of CV in an eternal AdS black hole background, since the volume grows at a monotonically increasing rate that approaches a constant from below at late times (e.g., see Fig. 7 in [257]). This constant exactly saturates Lloyd’s bound for large black holes. However, the rate of CA reaches a maximum at a finite time before approaching from above a constant final value saturating Lloyd’s bound (e.g., see Fig. 22 in [257]). Therefore, the intermediate regime in the evolution of CA violates the bound. Other counterexamples have also been found [258, 259, 260, 261, 262, 263, 264, 265, 266, 267]. It was argued in [268] that a black hole should be modeled by the composition of simple gates (i.e., close to the identity), and is thus incompatible with the hypotheses used to derive the Lloyd’s bound. Overall, the adaptation of Lloyd’s bound to holographic complexity is an open problem.
@@
Next, one can study the reaction of the holographic observables to shock waves corresponding to the insertion of null matter at arbitrary boundary times [253, 54]. Earlier, we assumed that the insertion times $t_{k}$ of the perturbations (which included the boundary times $t_{L},t_{R}$) were separated by an interval larger than $t_{*}$ (see the discussion below equation (122)). Here we will not make this assumption. For simplicity, let us fix $t_{L}=t_{R}=0$. We consider the case of a single time fold at the (negative) time $-t_{w}$ when the shock is inserted, and study the dependence of physical quantities on this time. In these cases, CV and CA show a plateau, where complexity is nearly constant, at early times $t_{w}\ll t_{*}$. But these quantities approach linear growth at later times $t_{w}\gg t_{*}$. The plateau becomes longer when the perturbation has smaller energy, showing signatures of scrambling in chaotic systems. [^29] Alternatively, one can study the full dependence of CV and CA on the boundary times, e.g., see Fig. 11 and Fig. 17 in Ref. [54]. In this case, holographic complexity admits a plateau region around $t_{L}=t_{R}=0$, which becomes longer when the shock is inserted at earlier times. This is another manifestation of the switchback effect. The scrambling time can be extracted by studying the plateau size, and turns out to be $t_{*}\approx\frac{1}{2\pi T}\log(M/T)$ (where $T$ is the Hawking temperature), as expected. Applying the CV conjecture, the results are summarized in Fig. 11. Furthermore, one can study the precise early $t_{w}$ dependence of holographic complexity. A careful analysis (summarized in Fig. 27 of [54]) shows that the complexity grows exponentially $\mathcal{C}(t_{w})\approx\exp(2\pi(t_{w}-t_{*})/\beta)$, with a characteristic exponent which is exactly equal to the Lyapunov exponent of maximally chaotic systems (see $\lambda_{\rm OTOC}$ in the key notion 6) [10]. Remarkably, all these behaviors of the quantities proposed as duals to holographic complexity agree with Eq. (11), obtained from the simple circuit models presented in Sec. 4.1. The matching is evident by comparing Figs. 11 and 5. Finally, [54] extends previous studies to cases where the shocks have energies above the thermal scale.
@@
Figure 11: Volume complexity $\Delta\mathcal{C}_{V}$ in three bulk (two boundary) dimensions, compared to the value $\Delta\mathcal{C}_{V,NS}$ of the complexity in the Neveu-Schwarz vacuum, as a function of the boundary insertion time $-t_{w}$ of the shock wave. $S_{1}$ is the thermal entropy of the black hole solution before the shock wave insertion; $T_{2}$ the Hawking temperature after the insertion of the shock. Picture taken from Fig. 19 of reference [54]. From bottom (green) to top (orange), the colors correspond to increasing value of the energy of the shockwave as follows: $E_{\text{shock}}/(2M_{1})=1.5\text{ (orange)},0.625\text{ (pink)},0.105\text{ (blue)},10^{-4}\text{ (red)},10^{-8}$ (green).
@@
[…]

### v3

@@ 7 Quantum complexity and space-time: a more concerted approach
Connections between geometry and entanglement were first proposed by Ryu and Takayangi Ryu:2006bv and were made covariant by Hubeny, Rangamani and Takayanagi Hubeny:2007xt. These authors showed that the entanglement entropies of subregions of the theory dual to an AdS spacetime are geometrized by the areas of certain extremal surfaces of the dual geometry. These developments led to major insights into the holographic duality, for example elucidating constraints on the entanglement patterns in quantum states that admit gravitational dual descriptions Hayden:2011ag. An analogy between the geometry of AdS spacetimes dual to conformal field theories and the graphical structure of tensor networks for constructing quantum states also provided an impetus, offering a schematic understanding of the hierarchical pattern of short and long range entanglement in holographic states Swingle:2009bg. Such developments motivated an important proposal by van Raamsdonk that spacetime is actually knitted together by the underlying patterns of quantum entanglement VanRaamsdonk:2010pw. Subsequent developments have illustrated many constraints on quantum states with holographic duals including new inequalities Hayden:2011ag and properties such a quantum error correction Almheiri:2014lwa (see Rangamani:2016dms for a review) that have a bearing on the complexity of constructing such states from a simple reference.
@@
Others proceeded to study the dynamics of thermalization via the AdS/CFT correspondence, especially considering the evolution of entanglement and the corresponding dual geometric quantities after quantum quenches Abajo-Arrastia:2010ajo, Balasubramanian:2011ur, Balasubramanian:2010ce, Balasubramanian:2011at, Liu:2013qca, Liu:2013iza. In all these cases, an initial rise of entanglement entropy ends in saturation after a thermalization time. These studies also led to an examination of the time evolution of entanglement entropy in thermo-field double (TFD) states dual to eternal black holes, i.e., to ERBs between two asymptotically AdS universes Hartman:2013aa, Maldacena:2013xja. The result, computed in terms of certain extremal surfaces passing through the ERB, shows that entanglement entropy grows with time in a TFD state until it saturates at equilibrium Hartman:2013aa. However, the black hole interior keeps growing long after the entanglement saturates. Therefore, Susskind suggested that the growth of the black hole’s interior should be encoded in the AdS/CFT correspondence by a dual quantum mechanical quantity that increases for a time exponential in the system’s size. What quantity could conceivably keep increasing after a system reaches thermal equilibrium? It is known in computer science that some computations carried out by the action of local operators can require exponentially long time or exponentially many actions of the operators, quantified as the “complexity” of the resulting state. If we think of time evolution as a computation performed by the action of a local Hamiltonian, one might guess that the complexity of the evolving quantum state, quantified in some way, may be related to the size of the black hole’s interior Susskind:2014rva.
@@
As we discussed in Secs. 4, 5, and 6, circuit complexity, Nielsen’s complexity, and Krylov/spread complexity all show a long period of linear growth for chaotic systems and also display the switchback effect (see discussion in Sec. 6.3.2). Therefore, any dual geometric object must also show these dynamics. Consistently with this, the initial conjecture was that the volume of the ERB wormhole is dual to some notion of time-evolving complexity of the dual Thermo-field Double state Susskind:2014rva. However, as we discuss in Sections 7.1 and 7.2, a number of different geometric quantities all reproduce linear growth and the switchback effect Stanford:2014jda, Brown:2015bva, Brown:2015lvg, Couch:2016exn, Belin:2021bga, Belin:2022xmt. We also discuss a proposal in the case of two-dimensional gravity for a non-perturbative definition of the size of the ERB wormhole Iliesiu:2021ari that saturates at times exponential in the black hole entropy, reproducing the expected late time saturation of complexity (Secs. 4,5 and 6). Strictly speaking, the various candidate holographic duals of complexity described in Sec. 7.1 all have divergences that require regulation. We understand these divergences as arising in the dual CFT from the continuum nature of the theory, just like the divergences of entanglement entropy. We explain these considerations in Sec. 7.3. Proposed extensions of the ideas to subregions, geometries with defects, and black holes in de Sitter space are summarized Sec. 7.4.
@@
All the geometric quantities proposed as duals to some notion of complexity in Sec. 7.1 are diffeomorphism-invariant, and hence we expect them to map onto precise gauge-invariant quantities in the dual theory. Is there a precise map to some particular notion of complexity? In other words, can we go beyond a qualitative agreement with the linear growth and switchback effect seen in any reasonable definition of complexity, to write an equation between gravitational and CFT quantities? Very recently, precise relations have been found between the size of an Einstein-Rosen Bridge in two-dimensional Jackiw-Teitelboim gravity and the spread complexity of a dual double-scaled SYK model, matching both the linear early time classical growth and the expected late time quantum saturation Berkooz:2018qkz, Lin:2019kpf, Rabinovici:2023yex, Nandy:2024zcd, Balasubramanian:2024lqk. There are also equalities relating the rate of change of Krylov/spread complexity and the momentum of particles in AdS spacetime. These results are reviewed in Sec. 7.5. Sec. 7.6 discusses some ideas for relating Nielsen’s complexity to dual geometries. Despite these findings, it is difficult to match geometrical observables in the gravitational setting with quantum complexity in field theory. This difficulty led to questions about the foundations of quantum computation in the context of the AdS/CFT correspondence. The existence of pseudorandom states, discussed in Sec. 7.7, can be used to argue that some entries in the holographic dictionary will be hard to compute, or alternatively, that quantum gravity efficiently solves problems that will require exponentially long times for a quantum computer. This unexpected line of research was revealed through the study of quantum complexity.
@@ 7.1 Holographic complexity proposals
Figure 10: Geometric objects defining holographic complexity proposals drawn inside the Penrose diagrams of the Schwarzschild-AdS black hole. Pictures adapted from Brown:2015lvg, Belin:2022xmt. (a) Maximal-volume codimension-one surface used in the CV proposal [Eq. (112)]. (b) WDW patch used to compute CA and CV2.0 in Eqs. (113) and (114). (c) Bulk region $\mathcal{M}$ with codimension-one boundaries $\Sigma_{\pm}$ defined by extremizing Eq. (115) in the CAny proposals.
The complexity equals volume (CV) proposal Stanford:2014jda, Susskind:2014moa declares that holographic complexity is proportional to the maximal volume $\mathcal{V}$ of a codimension-one spacelike surface $\mathcal{B}$ anchored at the (d1)(d-1)-dimensional boundary slice $\Sigma_{\rm CFT}$ where a CFT state is defined (see fig. 10(a)), i.e.,  [^30] In the context of CV conjecture, Ref. Couch:2018phr proposed to measure this geometric observable in terms of the maximal time from the horizon to the final slice (approached in the late stages of evolution). This notion also shares many similarities with the standard CV proposal, and at the same time removes certain ambiguities between small and large black holes in AdS spacetime.
@@
Other holographic proposals involve gravitational observables evaluated on the Wheeler-De Witt (WDW) patch, i.e., the bulk domain of dependence of a spacelike surface anchored at the boundary slice $\Sigma_{\rm CFT}$ (see fig. 10(b)). The first conjecture, complexity equals action (CA), associates complexity with the on-shell gravitational action $I_{\rm WDW}$ Brown:2015bva, Brown:2015lvg
@@
The gravitational action receives contributions from the null boundaries of the WDW patch, as required to make the variational principle in general relativity well-defined (see Lehner2016NullAct or appendix A of Carmi:2016wjl). The second proposal, complexity=volume 2.0 (CV2.0), identifies complexity as the spacetime volume of the WDW patch Couch:2016exn
@@
Each of these proposals presents ambiguities. CV and CV2.0 require a length scale $\ell_{\rm bulk}$ in the holographic dictionary. The CA proposal contains contributions from the null boundaries of the WDW patch, whose normal vectors depend on an arbitrary normalization constant. The last dependence can be removed by including certain counterterms in the gravitational action, but those in turn come equipped with their own (arbitrary) length scale, which we denote $L_{\rm ct}$ Lehner2016NullAct. When studying black holes, the CA and CV2.0 proposals probe regions near the black hole singularity. A delicate cancellation between a diverging metric factor and the vanishing volume of the sphere in the angular directions near the singularity leads to a result which does not diverge. Nevertheless, the validity of the formula in the semiclassical approximation is uncertain since the region near the singularity is not well understood in this limit.
The above ambiguities were taken as features of the holographic complexity proposal, rather than bugs, since computational complexity is also subject to several ambiguities: for instance, the choice of allowed gates in a circuit, the tolerance to reach a target state, the cost of gates etc., see Sec. 4. This perspective was taken a step further with a recent proposal that went under the name complexity equals anything (CAny), i.e., complexity may be dual to any one out of an infinite class of geometric observables. CAny requires a codimension-zero bulk region $\mathcal{M}$ with future (past) boundary $\Sigma_{+}(\Sigma_{-})$ anchored at the CFT slice $\partial\Sigma_{\pm}=\Sigma_{\rm CFT}$. One defines the following functional (see fig. 10(c)) Belin:2021bga, Belin:2022xmt, Jorstad:2023kmq
@@
where the shape of the boundaries $\Sigma_{\pm}$ is varied. When more than one solution exists, the one with maximal $W_{G_{2},F_{2,\pm}}$ is chosen. Once the unique codimension-zero region $(\bar{\mathcal{M}},\bar{\Sigma}_{\pm})$ solving Eq. (116) is identified, holographic complexity is computed by the observable
@@ 7.2 Time-dependent properties of the holographic proposals
$\mathcal{C}_{\rm Any}(t)=\frac{V_{x}}{G_{N}L}\sum_{\varepsilon=+,-}\int_{\Sigma_{\varepsilon}}d\sigma\,\mathcal{L}_{\varepsilon}(r,dr/d\sigma,dv/d\sigma)\,$ (120)
@@
Since $v$ is cyclic for $\mathcal{L}_{\varepsilon}$ in Eq. (120), [^31] This means that $\mathcal{L}_{\varepsilon}$ depends on $\dot{v}\coloneqq dv/d\sigma$, but not on $v$. the conjugate momenta $P_{v}^{\varepsilon}=\partial\mathcal{L}_{\varepsilon}/\partial\dot{v}$ are conserved along the profiles of $\Sigma_{\varepsilon}$. The late time regime $t\rightarrow\infty$ is achieved when each of the codimension-one surfaces $\Sigma_{\pm}$ hug a final slice at constant radial coordinate, such that the momenta approach constant values, up to exponentially-suppressed terms. This leads to the linear growth Belin:2021bga, Belin:2022xmt
@@
Let us consider the switchback effect, see e.g., Stanford:2014jda, Chapman:2018dem, Chapman:2018lsv. We modify the thermofield-double state as follows Shenker:2013pqa, Shenker:2013yza
@@
where $W_{L}$ are low-energy perturbations acting on the left boundary. These perturbations carry energies of the order of the black hole’s temperature (thermal-scale perturbations), which is assumed to be much smaller than the black hole’s mass. Let us assume that $|t_{k+1}-t_{k}|>t_{*}$, where $t_{*}$ is the scrambling time, and the label $k=0,1,\dots,n+1$ in the inequalities includes the boundary times via the definitions $t_{0}=-t_{R}$ and $t_{n+1}=t_{L}$. We also assume that, in the sequence of times $t_{0},\ldots t_{n+1}$ there are $n_{\rm sb}$ switchbacks (also called time folds): in other words, the arguments of the absolute values $|t_{k+1}-t_{k}|$ change sign $n_{\rm sb}\leq n$ times. To keep the bulk solutions as simple as possible, we assume that the perturbations are approximately spherically symmetric. The bulk geometry dual to the perturbed state (122) can be modeled by a black hole with $n\geq 1$ shock waves produced by null matter. In this setting, any CAny observable in the late time limit reads Belin:2022xmt
@@
The contribution $2n_{\rm sb}t_{*}$ reflects the spreading of the perturbations into the system during a time scale measured by the scrambling time $t_{*}=\frac{1}{2\pi T}\log\left(M/E\right)\approx\frac{1}{2\pi T}\log S$, where $E\ll M$ is the energy of the shock, and $S$ the black hole’s entropy. This characterizes the switchback effect for the CAny observables. At late boundary times $(t_{L},t_{R})$, holographic complexity will grow linearly with the sum of the boundary times $t_{L}+t_{R}$, but with a delay induced by the scrambling time of the perturbation. This phenomenon was already observed in earlier studies of the CV, CV2.0 and CA proposals Stanford:2014jda, Brown:2015lvg, Couch:2016exn. This is not a surprise. As we said, those proposals can be obtained from CAny by appropriate limiting procedures.
@@
Let us discuss other refined properties of the holographic conjectures, associated with early and intermediate times compared to the scrambling time and the inverse temperature scale. It was originally speculated that the black hole mass $M$ should provide an upper bound on the complexity rate $d\mathcal{C}/dt\leq 2M/\pi$, based on Lloyd’s bound on the computational speed Lloyd_2000. This statement was supported by the the full time evolution of CV in an eternal AdS black hole background, since the volume grows at a monotonically increasing rate that approaches a constant from below at late times (e.g., see Fig. 7 in Carmi:2017jqz). This constant exactly saturates Lloyd’s bound for large black holes. However, the rate of CA reaches a maximum at a finite time before approaching from above a constant final value saturating Lloyd’s bound (e.g., see Fig. 22 in Carmi:2017jqz). Therefore, the intermediate regime in the evolution of CA violates the bound. Other counterexamples have also been found Swingle:2017zcd, Yang:2017czx, Couch:2017yil, Auzzi:2018zdu, Auzzi:2018pbc, An:2018xhv, Alishahiha:2018tep, Bernamonti:2021jyu, Wang:2023ipy, Aguilar-Gutierrez:2023ccv. It was argued in Cottrell:2017ayj that a black hole should be modeled by the composition of simple gates (i.e., close to the identity), and is thus incompatible with the hypotheses used to derive the Lloyd’s bound. Overall, the adaptation of Lloyd’s bound to holographic complexity is an open problem.
@@
Next, one can study the reaction of the holographic observables to shock waves corresponding to the insertion of null matter at arbitrary boundary times Chapman:2018dem, Chapman:2018lsv. Earlier, we assumed that the insertion times $t_{k}$ of the perturbations (which included the boundary times $t_{L},t_{R}$) were separated by an interval larger than $t_{*}$ (see the discussion below equation (122)). Here we will not make this assumption. For simplicity, let us fix $t_{L}=t_{R}=0$. We consider the case of a single time fold at the (negative) time $t_{w}$ when the shock is inserted, and study the dependence of physical quantities on this time. In these cases, CV and CA show a plateau, where complexity is nearly constant, at early times $t_{w}\ll t_{*}$. But these quantities approach linear growth at later times $t_{w}\gg t_{*}$. The plateau becomes longer when the perturbation has smaller energy, showing signatures of scrambling in chaotic systems. [^32] Alternatively, one can study the full dependence of CV and CA on the boundary times, e.g., see Fig. 11 and Fig. 17 in Ref. Chapman:2018lsv. In this case, holographic complexity admits a plateau region around $t_{L}=t_{R}=0$, which becomes longer when the shock is inserted at earlier times. This is another manifestation of the switchback effect. The scrambling time can be extracted by studying the plateau size, and turns out to be $t_{*}\approx\frac{1}{2\pi T}\log(M/T)$ (where $T$ is the Hawking temperature), as expected. Applying the CV conjecture, the results are summarized in Fig. 11. Furthermore, one can study the precise early $t_{w}$ dependence of holographic complexity. A careful analysis (summarized in Fig. 27 of Chapman:2018lsv) shows that the complexity grows exponentially $\mathcal{C}(t_{w})\approx\exp(2\pi(t_{w}-t_{*})/\beta)$, with a characteristic exponent which is exactly equal to the Lyapunov exponent of maximally chaotic systems (see $\lambda_{\rm OTOC}$ in the key notion 6) Maldacena:2015waa. Remarkably, all these behaviors of the quantities proposed as duals to holographic complexity agree with Eq. (11), obtained from the simple circuit models presented in Sec. 4.1. The matching is evident by comparing Figs. 11 and 5. Finally, Chapman:2018lsv extends previous studies to cases where the shocks have energies above the thermal scale.
@@
Figure 11: Volume complexity $\Delta\mathcal{C}_{V}$ in three bulk (two boundary) dimensions, compared to the value $\Delta\mathcal{C}_{V,NS}$ of the complexity in the Neveu-Schwarz vacuum, as a function of the boundary insertion time $t_{w}$ of the shock wave. $S_{1}$ is the thermal entropy of the black hole solution before the shock wave insertion; $T_{2}$ the Hawking temperature after the insertion of the shock. Picture taken from Fig. 19 of reference Chapman:2018lsv. From bottom (green) to top (orange), the colors correspond to increasing value of the energy of the shockwave as follows: $E_{\text{shock}}/(2M_{1})=1.5\text{ (orange)},0.625\text{ (pink)},0.105\text{ (blue)},10^{-4}\text{ (red)},10^{-8}$ (green).
@@
In the case of the CV conjecture, the answer to both questions is affirmative for two-dimensional models of dilaton gravity, including JT gravity. Reference Iliesiu:2021ari proposed a nonperturbative definition of the ERB’s length that includes quantum corrections from surfaces with higher topologies. We denote by $\gamma$ any non self-intersecting geodesic with length $\ell_{\gamma}$, $\Delta>0$ a regulator, and $\langle\dots\rangle$ the sum over geodesics defined over surfaces of any topology within the gravitational path integral. The nonperturbative length $\ell$ of the ERB reads